A line passes through the orgin and has a slpoe of -3. what is the equation of the line that is perpendicular to the first line and passes through the point (3,4)?
Answer by stanbon(75887)

A line passes through the orgin and has a slpoe of -3. what is the equation of the line that is perpendicular to the first line and passes through the point (3,4)?
----
The slope of the line you want must be 1/3
-----
Find equation of line with m = 1/3 and y = 4 when x = 3.
------
Equation Form:
y = mx + b
4 = (1/3)(3) = b
b = 4-1
b = 3
------
Equation: Y = (1/3)x+3
==========================
Cheers,
Stan H.
==========================